OpenStudy (anonymous):

One solution of a quadratic equation with real coefficients is m-i√(n), find the equation

OpenStudy (anonymous):

your job is to multiply out \[(x-(n+\sqrt{n}i)(x-(m-\sqrt{n}i))\]which is not has hard as it seems.

OpenStudy (anonymous):

the constant will be \[m^2+n\] and the "middle term" will be \[-2mx\] so your equation will be \[x^2-2mx+m^2+n=0\]

OpenStudy (anonymous):

typo on first line should be \[(x-(m+\sqrt{n}i)(x-(m-\sqrt{n}i))\]

OpenStudy (anonymous):

how do you find the constant?

OpenStudy (anonymous):

when you multiply a complex number \[a+bi\] by its complex conjugate \[a-bi\] you get \[a^2+b^2\]

OpenStudy (anonymous):

in this case you have \[(m+\sqrt{n}i)(m-\sqrt{n}i)=m^2+n\]

OpenStudy (anonymous):

thanks